The director of research and development is testing a new medicine. She wants to know if there is evidence at the 0.02 level that the medicine relieves pain in more than 349 seconds. For a sample of 99 patients, the mean time in which the medicine relieved pain was 353 seconds. Assume the population standard deviation is 19. Find the P-value of the test statistic.
Answer #1

Null Hypothesis H0 : μ = 349

Alternative Hypothesis H1:u> 349

Under H0, the test statistic is

7-μ = 353-349 = 2.095

The critical value of Z at 2% level of significance is 2.05

The P-Value is .0181

Since Z calculated is greater than Z tabulated, Reject H0.

Hence we have enough evidence to support the claim that the medicine relieves pain in more than 349 seconds.